# Vision Transformers in Precision Agriculture: A Comprehensive Survey

### v3

@@ 4.2 Hybrid Models
In another study [98], a parallel fusion of CNN and ViT modules was proposed to capture both local and global features. Unlike sequential CNN-ViT hybrids, this approach enables independent feature extraction followed by fusion using MobileNetV2 [79] and Squeeze-and-Excitation (SE) blocks, thereby improving efficiency and interpretability. ConViTX is ultra-lightweight, comprising only 704,882 trainable parameters and requiring 0.647 GFLOPs, making it ideal for deployment on resource-constrained IoT/edge devices. Explainability is achieved using Grad-CAM and LIME, and the model’s feature discrimination capability is further validated through t-SNE visualizations. The model was rigorously evaluated on five datasets: PlantVillage, Embrapa, PlantDoc, a custom in-field maize dataset (IIITDM-Maize), and a combined multi-species dataset called PlantCOMB. On PlantVillage, ConViTX achieved an accuracy of 99.63%, an F1-score of 99.64%, and an AUC of 99.96%, outperforming nine state-of-the-art (SOTA) models. On Embrapa, it achieved an F1-score of 95.12% and an AUC of 99.74%, and on PlantDoc, it retained the top spot in most metrics. On the custom maize dataset, it achieved 93.02% accuracy, matching the best competitors while showing a 2% higher AUC. On raw drone images, it reached 61.42% accuracy, surpassing most SOTA models in generalizability. Beyond static evaluation, ConViTX was deployed on a Jetson Nano-equipped drone and tested on real-time video frames captured from maize farms. It processed frames at approximately 29 FPS and demonstrated competitive results, notably identifying northern leaf blight with 58% accuracy and healthy plants with 73% accuracy. Though challenges remain with subtle diseases like common rust, the model achieved a better balance between speed, accuracy, and interpretability compared to other lightweight and heavyweight models. Overall, ConViTX proves to be an efficient, interpretable, and deployable solution, opening a promising avenue for hybrid CNN-ViT architectures in real-world smart agriculture systems.
RepAgrViT [18], a novel lightweight hybrid architecture, is proposed to integrate CNN and ViT principles for plant disease recognition in Internet of Things (IoT) edge environments. Unlike traditional CNN-ViT hybrids, RepAgrViT employs a unique dual-stream design consisting of a token mixer and a channel mixer arranged in series to simultaneously process local and global features. A core innovation is the Bilinear Attention Transformation (BATransform) module, which captures long-range dependencies between diseased and healthy leaf regions. Additionally, the model integrates structural reparameterization techniques to reduce inference complexity while preserving accuracy. At just 9.6M parameters, the model demonstrates strong computational efficiency while retaining competitive performance.
To optimize hyperparameters efficiently, the authors introduce Lite-AVPSO, a novel Particle Swarm Optimization variant that incorporates adaptive weighted delayed velocity and neighborhood-based local search mechanisms. This approach significantly enhances the hyperparameter search process over alternatives like TPE, GA, and Hyperband. Through an extensive optimization process spanning over 200 trials, the model achieved 98.54% average accuracy across three plant disease datasets: the Coffee Leaf Dataset , the Plant Pathology Dataset , and the Rice Leaf Disease Dataset. On these datasets, RepAgrViT achieved 98.03%, 96.99%, and 98.78% accuracy respectively. Overall, this work contributes a highly efficient and interpretable AI system tailored for precision agriculture in resource-constrained settings.
@@ 5.2 Open Challenges
Another challenge is the high computational cost and resource demands of ViTs. The self-attention mechanism that enables ViTs to capture long-range dependencies within images is computationally expensive, particularly when dealing with the high-resolution images common in agricultural applications. This leads to high memory and processing power requirements, making it difficult to deploy ViTs on resource-limited devices such as drones or edge sensors used in agricultural fields. Additionally, training ViTs can take considerably longer than training CNN-based models, which may hinder their use in real-time agricultural applications where fast responses are essential. Lastly, generalizing across varied agricultural environments presents another significant challenge. While ViTs have demonstrated strong performance on specific datasets, they may not be as effective in different climates, crop types, or farming practices. The agricultural landscape is highly diverse, and models trained on data from one region or crop type may fail to adapt to new environments with different lighting conditions, backgrounds, or plant variations. This lack of environmental robustness could limit the practical application of ViTs in precision agriculture on a global scale. Therefore, enhancing the ability of ViTs to generalize across diverse conditions—through methods like domain adaptation or synthetic data generation—is a critical area for future research. We can summarize the open challenges as:
@@
PlantVillage [60] Contains over 50,000 images of plant leaves across 14 species and 26 diseases $20$
@@
PlantPathology2020 [99] Contains over 3,000 high-quality real-life images of apple leaf diseases under varying conditions $4$
PlantDoc [89] Contains over 2,000 images across 13 plant species and 17 disease classes $4$
CD&S [3] Contains over 4,000 corn images comprising field images and augmented images $2$
Embrapa [5] Contains over 56,000 images of 171 diseases and other disorders affecting 21 plant species $2$
PaddyDoctor [67] Contains over 16,000 annotated paddy leaf images across 13 classes $1$
Plant Seedling [27] Contains over 900 unique plants belonging to 12 species $1$
PlantCLEF2022 [43] Contains over 2,000,000 images and 80,000 classes $1$
DRLI [16] Contains over 4,000 images of healthy and diseased leaves $1$
PlantComposite [65] Contains over 58,000 images of healthy and diseased leaves $1$
Citrus Fruits and Leaves dataset [71] Contains over 600 images of citrus leaves and 150 images of citrus fruits, with five classes for each $1$
Dhan-Shomadhan [36] Contains over 1,100 images of five rice leaf diseases $1$
Strawberry Disease dataset [1] Contains 2,500 images of seven types of strawberry diseases $1$
Maize and Rice dataset [15] Contains 500 rice images and 466 maize images $1$
@@ 6 Conclusion
The use of ViTs in agriculture marks a noteworthy deviation from traditional DL paradigms, with promising potential to model complex visual patterns without relying on the strong inductive biases of convolutional architectures. While CNNs have traditionally excelled at agricultural tasks due to their locality and translation equivariance, ViTs offer a more versatile and scalable approach, particularly when combined with appropriate training methods and hybrid models that reintroduce inductive biases in a controlled manne
In this survey, we have highlighted the evolution of ViT-based approaches in agriculture, ranging from vanilla Transformer models to hybrid models that leverage the strengths of both CNNs and Transformers. The literature presented demonstrates the potential of ViTs for a range of agricultural tasks such as disease identification, yield prediction, and precision agriculture, with most models outperforming or complementing traditional approaches. Despite these advances, several open issues remain, including data scarcity, computational demands, and the need for task-specific model customization. Addressing these challenges will be essential to fully realizing the potential of ViTs in real-world agricultural applications. As the research field continues to mature, future work must prioritize model efficiency, domain adaptation, and the creation of robust, annotated agricultural datasets. In conclusion, ViTs represent a compelling paradigm shift for agricultural vision tasks. As innovation and synergy between agricultural sciences and machine learning continue, they are likely to drive smarter, more sustainable, and more precise agricultural solutions in the future.

### v4

@@ 4.2 Hybrid Models
A real-time plant disease identification framework called CondConViT [98] has been proposed, which integrates a ViT with a Conditional Convolutional Neural Network (CondConv) and is further enhanced by a novel attention mechanism known as Conditional Attention with Statistical Squeeze-and-Excitation (CASSE). To improve generalization and robustness, the authors introduced a data augmentation technique based on a modified CycleGAN, termed C3GAN, which synthesized realistic in-field images from healthy plant samples. The model was designed to be lightweight, with only 0.95 million parameters, and was deployable on edge devices for drone-based surveillance. CondConViT was evaluated against seven state-of-the-art models and demonstrated superior interpretability and classification performance across diverse environmental conditions using both Grad-CAM and LIME techniques. They used six datasets, including five public datasets: PlantVillage [61], Embrapa [5], Plantpathology [100], Maize and Rice [15], and a newly created in-field drone image dataset named IIITDMJ-Maize [97], which comprised 416 labeled images of maize plants captured under various weather conditions. C3GAN was employed to augment the IIITDMJ-Maize dataset by generating 300 synthetic images for three maize diseases: common rust, northern leaf blight, and gray leaf spot. The CondConViT model achieved the highest or near-highest performance across all datasets in terms of accuracy, F1-score, and AUC. Furthermore, it exhibited strong generalization on unseen raw drone images and outperformed heavier models, making it suitable for real-time and resource-constrained agricultural applications. In another study [99], a parallel fusion of CNN and ViT modules was proposed to capture both local and global features. Unlike sequential CNN-ViT hybrids, this approach enables independent feature extraction followed by fusion using MobileNetV2 [80] and Squeeze-and-Excitation (SE) blocks, thereby improving efficiency and interpretability. ConViTX is ultra-lightweight, comprising only 704,882 trainable parameters and requiring 0.647 GFLOPs, making it ideal for deployment on resource-constrained IoT/edge devices. Explainability was achieved using Grad-CAM and LIME, and the model’s feature discrimination capability was further validated through t-SNE visualizations. The model was rigorously evaluated on five datasets: PlantVillage, Embrapa, PlantDoc, a custom in-field maize dataset (IIITDM-Maize), and a combined multi-species dataset called PlantCOMB. On PlantVillage, ConViTX achieved an accuracy of 99.63%, an F1-score of 99.64%, and an AUC of 99.96%, outperforming nine state-of-the-art (SOTA) models. On Embrapa, it achieved an F1-score of 95.12% and an AUC of 99.74%, and on PlantDoc, it retained the top spot in most metrics. On the custom maize dataset, it achieved 93.02% accuracy, matching the best competitors while showing a 2% higher AUC. On raw drone images, it reached 61.42% accuracy, surpassing most SOTA models in generalizability. Beyond static evaluation, ConViTX was deployed on a Jetson Nano-equipped drone and tested on real-time video frames captured from maize farms. It processed frames at approximately 29 FPS and demonstrated competitive results, notably identifying northern leaf blight with 58% accuracy and healthy plants with 73% accuracy. Though challenges remain with subtle diseases like common rust, the model achieved a better balance between speed, accuracy, and interpretability compared to other lightweight and heavyweight models. Overall, ConViTX proves to be an efficient, interpretable, and deployable solution, opening a promising avenue for hybrid CNN-ViT architectures in real-world smart agriculture systems. RepAgrViT [18], a novel lightweight hybrid architecture, was proposed to integrate CNN and ViT principles for plant disease recognition in Internet of Things (IoT) edge environments. Unlike traditional CNN-ViT hybrids, RepAgrViT employs a unique dual-stream design consisting of a token mixer and a channel mixer arranged in series to simultaneously process local and global features. A core innovation is the Bilinear Attention Transformation (BATransform) module, which captures long-range dependencies between diseased and healthy leaf regions. Additionally, the model integrates structural reparameterization techniques to reduce inference complexity while preserving accuracy. At just 9.6M parameters, the model demonstrates strong computational efficiency while retaining competitive performance. To optimize hyperparameters efficiently, the authors introduce Lite-AVPSO, a novel Particle Swarm Optimization variant that incorporates adaptive weighted delayed velocity and neighborhood-based local search mechanisms. This approach significantly enhances the hyperparameter search process over alternatives like TPE, GA, and Hyperband. Through an extensive optimization process spanning over 200 trials, the model achieved 98.54% average accuracy across three plant disease datasets: the Coffee Leaf Dataset , the Plant Pathology Dataset , and the Rice Leaf Disease Dataset. On these datasets, RepAgrViT achieved 98.03%, 96.99%, and 98.78% accuracy respectively. Overall, this work contributes a highly efficient and interpretable AI system tailored for precision agriculture in resource-constrained settings.
@@ 5.2 Open Challenges
Another challenge is the high computational cost and resource demands of ViTs. The self-attention mechanism that enables ViTs to capture long-range dependencies within images is computationally expensive, particularly when dealing with the high-resolution images common in agricultural applications. This leads to high memory and processing power requirements, making it difficult to deploy ViTs on resource-limited devices such as drones or edge sensors used in agricultural fields. Additionally, training ViTs can take considerably longer than training CNN-based models, which may hinder their use in real-time agricultural applications where fast responses are essential. Lastly, generalizing across varied agricultural environments presents another significant challenge. While ViTs have demonstrated strong performance on specific datasets, they may not be as effective in different climates, crop types, or farming practices. The agricultural landscape is highly diverse, and models trained on data from one region or crop type may fail to adapt to new environments with different lighting conditions, backgrounds, or plant variations. This lack of environmental robustness could limit the practical application of ViTs in precision agriculture on a global scale. Therefore, enhancing the ability of ViTs to generalize across diverse conditions is a critical area for future research. We can summarize the open challenges as:
@@
PlantVillage [61] Contains over 50,000 images of plant leaves across 14 species and 26 diseases 2020
@@
PlantPathology2020 [100] Contains over 3,000 high-quality real-life images of apple leaf diseases under varying conditions 44
PlantDoc [90] Contains over 2,000 images across 13 plant species and 17 disease classes 44
CD&S [2] Contains over 4,000 corn images comprising field images and augmented images 22
Embrapa [5] Contains over 56,000 images of 171 diseases and other disorders affecting 21 plant species 22
PaddyDoctor [68] Contains over 16,000 annotated paddy leaf images across 13 classes 11
Plant Seedling [27] Contains over 900 unique plants belonging to 12 species 11
PlantCLEF2022 [44] Contains over 2,000,000 images and 80,000 classes 11
DRLI [16] Contains over 4,000 images of healthy and diseased leaves 11
PlantComposite [66] Contains over 58,000 images of healthy and diseased leaves 11
Citrus Fruits and Leaves dataset [72] Contains over 600 images of citrus leaves and 150 images of citrus fruits, with five classes for each 11
Dhan-Shomadhan [37] Contains over 1,100 images of five rice leaf diseases 11
Strawberry Disease dataset [1] Contains 2,500 images of seven types of strawberry diseases 11
Maize and Rice dataset [15] Contains 500 rice images and 466 maize images 11
@@ 6 Conclusion
The adoption of ViTs in agriculture represents a significant departure from conventional DL approaches, offering considerable potential for capturing complex visual patterns without relying on the strong inductive biases inherent in convolutional architectures. While CNNs have historically demonstrated exceptional performance in agricultural applications due to their locality and translation equivariance, ViTs present a more adaptable and scalable alternative. Their effectiveness is further enhanced when combined with specialized training strategies and hybrid architectures that strategically reintroduce inductive biases, enabling a balanced trade-off between flexibility and domain-specific robustness. Collectively, these advancements indicate that ViTs, whether as standalone models or within hybrid frameworks, are poised to play a pivotal role in the future of agricultural applications.
In this survey, we have traced the evolution of ViT-based approaches in agriculture, from vanilla Transformer models to hybrid architectures that leverage the strengths of both CNNs and Transformers. The reviewed literature demonstrates the potential of ViTs across a range of agricultural tasks, including disease identification, yield prediction, and precision agriculture, with most models outperforming or complementing traditional methods. Despite these advances, several challenges remain, including data scarcity, high computational requirements, and the need for task-specific model customization. Addressing these challenges is crucial for fully realizing the potential of ViTs in practical agricultural applications. As the field progresses, future research should prioritize the enhancement of model efficiency, the adaptation of models to diverse agricultural domains, and the creation of robust, meticulously annotated agricultural datasets.